OpenStudy (anonymous):

Solve the system using any method. 5x-5y=0 7x-8y=8

OpenStudy (freckles):

notice the first equation can be written as 5x=5y or x=y just sub this into the second equation

OpenStudy (anonymous):

 Is this how you set it up?

OpenStudy (abhisar):

Solve the system using any method. 5x-5y=0 7x-8y=8 You can write the first equation as \(8y-8x=0\) Add this to the second equation and you will an equation in terms of x only. Find the value of x and then substitute it in any one of the parent equation to get value of y

OpenStudy (anonymous):

why would the 1st equation into 8y=8x=0?

OpenStudy (abhisar):

\(5x-5y=0 \\ \Rightarrow 5(x-y)=0 \\ \Rightarrow 8x-8y=0 \Rightarrow 8y-8x=0\)

OpenStudy (abhisar):

Right? Now the second equation has -8y, so we can cancel them out out and life will become happy c:

OpenStudy (midhun.madhu1987):



OpenStudy (abhisar):

Yes, you can do that way too

OpenStudy (anonymous):

Oh i understand now. :D Thanks so much!